OpenStudy (anonymous):

solve by completing the square. x2^2-4x-1=0

OpenStudy (anonymous):

is it supposed to be 2x^(2)-4x-1=0 ?

OpenStudy (solomonzelman):

@sierra1003 ? yes or no ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I sorry for the late respond I have a ton of homework and little time to do it.

OpenStudy (solomonzelman):

you are trying to make the left side \[x^2+2x+1~~~~b/c~~i t's~~perfect~~squ are\]

OpenStudy (solomonzelman):

add 3 to both sides \[ 2x^2-4x-1=0\]\[ 2x^2-4x-1+3=0+3\]\[ 2x^2-4x+2=3\]\[2(x^2+2x+1)=3\]\[x^2+2x+1=3/2\]\[(x+1)^2=3/2\]

OpenStudy (solomonzelman):

\[(x+1)^2= \frac{3}{2}\]\[x+1=±\sqrt{ \frac{3}{2} }\]\[x+1=± \frac{ \sqrt{3 } }{ \sqrt{2 } } \]\[x+1=± \frac{ \sqrt{6 } }{ 2 } \]\[x=-1± \frac{ \sqrt{6 } }{ 2 } \]\[x=-\frac{2}{2}± \frac{ \sqrt{6 } }{ 2 } \]\[x= \frac{ -2±\sqrt{6 } }{ 2 } \]see what I am doing ?